OpenStudy (anonymous):

Use the distributive property to simplify the expression below: -5x-15/-5 −x + 3 x − 15 x − 3 x + 3

OpenStudy (anonymous):

I'm Thinking It's "A". -_-

OpenStudy (anonymous):

-5x/-5=x -15/-5=3 x+3

OpenStudy (anonymous):

Ok. Thanks! @Grazes

OpenStudy (anonymous):

No probs. But you know how I got that, right? Just separate the expression?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

You can alternatively do this if you want to :)\[\frac{-5x - 15}{-5} = \frac{-5(x + 3)}{-5} = \frac{\cancel{-5}(x + 3)}{\cancel{-5}} = x + 3\]